### v2

# Quantitative imaging of nonlinear spin-wave propagation using diamond quantum sensors

Abstract
Spin waves propagating in magnetic materials exhibit nonlinear behavior at large amplitudes due to the competition between excitation and relaxation, providing an attractive platform for exploring nonlinear wave dynamics. In particular, spin waves with a non-zero wavenumber that carry momentum undergo nonlinear relaxation and experience wavenumber modulation in the nonlinear regime. This nonlinearity has been observed experimentally—for example, in S. R. Lake et al., Phys. Rev. Appl. 17, 034010 (2022)— but a quantitative comparison with theory has not yet been carried out. Here, We image nonlinear spin-wave propagation in two yttrium iron garnet thin films with distinct spin-wave decay rates using a wide-field quantum diamond microscope. We obtain quantitative distributions of spin-wave amplitude and phase as a function of the excitation microwave strength. As a result, we observe a threshold in the spin-wave amplitude beyond which nonlinear effects become evident and confirm that this threshold is consistent with theoretical predictions based on four-magnon scattering for both samples. Moreover, as the amplitude of the spin waves increases, we observe modulation of the wavenumber across the field of view. We attribute this modulation primarily to a reduction in the saturation magnetization caused by incoherent spin waves generated by multi-magnon scattering. Our quantitative measurements provide a pathway for visualizing nonlinear spin-wave dynamics and are crucial for deepening our understanding of the underlying mechanisms.
December 11, 2025
I Introduction
Spin waves, collective excitations of ordered magnets, possess an intrinsic nonlinearity [1, 2]. This nonlinearity arises from the competition between spin-wave excitation and relaxation and becomes prominent when the spin-wave amplitude exceeds a certain threshold. The nonlinear dynamics of the uniform mode (i.e., ferromagnetic resonance, FMR) have been comprehensively studied since the 1950s. In these studies, various nonlinear processes have been explored, including parametric excitation [3, 4, 5], Bose-Einstein condensation of magnons [6], and the first and second Suhl instabilities [1]. Recently, significant interest has focused on the dynamics of spin waves with non-zero wavenumbers due to their rich physics and potential applications in information devices [7]. The nonlinear behavior of these spin waves has been investigated both theoretically and experimentally [8, 9, 10, 11, 12, 13], in which the initially excited coherent spin waves are scattered by multi-magnon processes, thereby magnons with various wavenumbers and frequencies are generated. In addition, when the spin-wave amplitude becomes large, an effective reduction in the static magnetization leads to a frequency shift of the spin waves [2, 13, 14].
These nonlinear dynamics of the spin waves are primarily investigated using a vector network analyzer (VNA) [12], magneto-optical effects [15], and, especially in recent years, Brillouin light scattering (BLS) [8, 9, 10, 13]. BLS enables time-, space-, and frequency-resolved detection of spin-wave propagation, making it a valuable technique for visualizing these dynamics [16]. However, the signal obtained via BLS is the intensity of the scattered light arising from Raman scattering processes. While this intensity is proportional to the magnon intensity, there is no direct theoretical relationship that links it to the spin-wave amplitude. Although Ref. [13] systematically investigates the nonlinearity of surface spin waves propagating on a magnetic waveguide, discusses the underlying mechanism, and estimates the precession angle at which nonlinearity emerges by calibrating the spin-wave amplitude from the variation in wavenumber, quantitative comparisons with theoretical predictions have not been carried out, and thus the validity of the obtained amplitude has not yet been verified.
Recently, nitrogen-vacancy (NV) centers in diamonds [17] have gained significant attention as quantitative sensors for microwave amplitude. With a resonant frequency of approximately $3\,\mathrm{GHz}$, they are particularly sensitive to microwave fields in this frequency range, making them well suited for sensing spin waves, whose frequencies typically lie in the several GHz range. In previous studies, it has been demonstrated that both the amplitude and phase of coherently propagating spin waves in magnetic films can be quantitatively imaged [18, 19, 20, 21, 22].
Although some previous studies employing NV centers has observed nonlinear spin-wave dynamics, such as the generation of incoherent spin waves at the resonant frequencies of the NV spins when exciting the FMR mode [23, 24] and the frequency conversion of coherent spin waves by spin-wave mixing through four-magnon scattering [25], no study to date has quantitatively examined the nonlinear dynamics of propagating spin waves with non-zero wavenumbers themselves.
Figure 1: (a) Schematic of the experiment. (b) Dispersion relation of the surface spin waves for YIG-A at an external magnetic field $B_{0}=25.17\,\mathrm{mT}$. The horizontal dashed line corresponds to the resonance frequency of the NV spin. (c) Two-dimensional image of the microwave amplitude at an external magnetic field $B_{0}=25.17\,\mathrm{mT}$ and input microwave power $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ for YIG-A. (d) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for the data in (c).
In this paper, we present results on imaging the amplitude and phase of surface spin waves in two yttrium iron garnet (YIG) thin films with distinct characteristics using a wide-field quantum diamond microscope (QDM), while systematically varying the excitation microwave amplitude. First, we show that the spin waves in the first YIG thin film, grown by sputtering, exhibit a significant decay beyond a characteristic amplitude threshold. We confirm that this threshold is consistent with theoretical predictions based on four-magnon scattering. In addition, as the amplitude of the spin waves increases, we observe modulation of the wavenumber within the field of view, and our quantitative analysis reveals that this modulation largely originates from incoherent spin waves generated by multi-magnon scattering. Finally, to validate our findings, we show that the second YIG thin film, which is grown by liquid phase epitaxy (LPE) and characterized by a smaller relaxation rate than the first YIG film, exhibits a distinct spin-wave amplitude threshold, also in agreement with theoretical expectations. This achievement demonstrates the usefulness of quantitative QDM investigations for probing nonlinear spin-wave dynamics and contributes to a systematic understanding of nonlinear spin waves.
This paper is organized as follows. We describe the experimental setup and principle in Sec. II. In Sec. III.1, we show spin-wave imaging in the linear regime for the first YIG film. We then present the nonlinear behavior that emerges as the spin-wave amplitude is increased in Sec. III.2, and quantitatively extract the spin-wave amplitude in Sec. III.3. Then, we introduce the theoretical formulation to discuss the threshold between the linear and nonlinear regimes and apply it to our experimental findings in Sec. III.4. We address the modulation of the wavenumber in Sec. III.5. Sec. III.6 is devoted to comparing with the second YIG film. Finally, we conclude the paper with a summary and outlook in Sec. IV.
II Methods
The experimental setup is depicted in Fig. 1(a), where we employ a QDM as the measurement system. In this setup, a green laser (wavelength: $515\,\mathrm{nm}$) irradiates an ensemble of NV centers, and their red fluorescence is collected using an sCMOS camera [26]. The magnification of the microscope is $100\times$, and the field of view is approximately $100\,\mathrm{\mu m}$ square.
The diamond chip used in this study is fabricated by cutting a (111) oriented Ib-type diamond single crystal (Element Six, MD111) into a substrate with a thickness of $38\,\mathrm{\mu m}$ in the (110) orientation. An NV ensemble layer is formed by carbon ion implantation ($30\,\mathrm{keV}$, $1\times 10^{12}\,\mathrm{cm^{-2}}$) and subsequent anneals [27, 26]. This process creates an NV layer at a depth of around $30\,\mathrm{nm}$ from the surface. We utilize NV centers oriented in the in-plane direction [along the $y$-axis; see Fig. 1(a)].
We use two YIG thin films with different growth methods as our magnetic materials. The first film, “YIG-A”, is epitaxially grown on (111)-oriented gadolinium gallium garnet (GGG) substrate using RF sputtering in the group of the co-authors of this paper [28], and has a thickness of $d=54\,\mathrm{nm}$. The second film, “YIG-B”, is epitaxially grown on a (111)-oriented GGG substrate by LPE and is commercially obtained from Matesy GmbH, with a thickness of $d=109\,\mathrm{nm}$. For YIG-A, the measurement setup is identical to that described in Ref. [26] except for the configuration of the microwave circuits.
Striplines with a width of approximately $20\,\mathrm{\mu m}$ and a length of $2\,\mathrm{mm}$, with $500\,\mathrm{\mu m}\times 500\,\mathrm{\mu m}$ pads for wiring at both ends, are deposited on the film surface for both YIG samples. Microwaves are input to the stripline on the left side to excite spin waves [Fig. 1(a)]. The striplines on YIG-A consist of Ti/Cu/Au ($10\,\mathrm{nm}/190\,\mathrm{nm}/9\,\mathrm{nm}$), and those on YIG-B consist of Ti/Cu/Au ($10\,\mathrm{nm}/97\,\mathrm{nm}/9\,\mathrm{nm}$). Microwaves are first amplified from the signal generator through an amplifier and are then fed into the coplanar waveguide; from there, they are delivered to the stripline by wire bonding gold wires to the pad. The other side is also connected to the coplanar waveguide and terminated with $50\,\Omega$. We define the input microwave power $P_{\mathrm{mw}}$ as the power supplied to the coplanar waveguide. Consequently, due to differences in losses, the actual microwave power reaching the stripline may differ between YIG-A and YIG-B.
The diamond chips are attached to the YIG films using varnish, with the surface containing the NV layer facing downward. The distance $z_{d}$ between the NV layer and the surface of the magnetic film is estimated by evaluating the distribution of the stray magnetic field from the stripline on the surface of the YIG film under current application [29, 26]. For YIG-A and YIG-B, the distance $z_{d}$ is estimated to be $(878\pm 20)\,\mathrm{nm}$ and $z_{d}=(860\pm 18)\,\mathrm{nm}$, respectively (see Supplemental Materials for details [30]).
A static magnetic field is applied in the film plane along the stripline ($y$-axis) using a Helmholtz coil. In this configuration, surface spin waves propagating perpendicular to the magnetic field are excited [31].
Using Rabi oscillation measurements, we detect spin waves whose frequency matches the resonance between the $m_{s}=0$ and $m_{s}=-1$ states of the NV spins [32, 19, 21]. In each region of interest, we fit the Rabi oscillation signals and obtain the Rabi frequency, which is then converted to the amplitude of the circularly polarized microwave field $B_{\mathrm{mw}}$ [26].
III Results and Discussions
III.1 Spin-wave imaging in the linear regime
First, we present the results for YIG-A. The external magnetic field is set to $B_{0}=25.17\,\mathrm{mT}$. Figure 1(b) shows the dispersion relation of the surface spin waves under this magnetic field as well as the resonance frequency of the NV spins. The dispersion relation is given by the following equation [33, 34, 35]:
$\begin{split}f_{\mathrm{sw}}(k_{x})=\gamma_{e}\mu_{0}&\sqrt{\quantity(H_{0}+M_{\mathrm{eff}}(1-g(k_{x}d)))\quantity(H_{0}+M_{\mathrm{eff}}g(k_{x}d))},\\ g(k_{x}d)&=1-\frac{1-e^{-k_{x}d}}{k_{x}d},\end{split}$ (1)
where $\gamma_{e}$ is the gyromagnetic ratio of an electron spin, $\mu_{0}$ is the permeability of vacuum, $M_{\mathrm{eff}}\,(=169.6\,\mathrm{mT})$ is the effective magnetization [26], and $d=54\,\mathrm{nm}$ is the thickness of the YIG film. Under this magnetic field, the resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2165.8\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.4\,\mathrm{rad/\mu m}$.
Figure 1(c) shows the two-dimensional distribution of the microwave field amplitude at an input microwave power of $P_{\mathrm{mw}}=11\,\mathrm{dBm}$. In the analysis, signals from each $520\,\mathrm{nm}$ in the $x$- and $y$-directions are integrated and treated as a single pixel. The microwave amplitude oscillates periodically in the $x$-direction due to the interference between the microwave field generated by the spin waves propagating in the $x$-direction and the reference microwave with a spatially uniform phase across the entire field of view [21, 22, 26]. In addition to the propagation in the $x$-direction, an increase in the microwave amplitude is observed at a specific diagonal angle. This phenomenon, known as “caustics,” arises from the anisotropy of the spin-wave dispersion relation [36, 21, 26].
Figure 1(d) displays the one-dimensional microwave amplitude distribution along the $x$-direction $B_{\mathrm{mw}}(x)$. Here, signals are integrated over a width of $13\,\mathrm{\mu m}$ from $y=31.87\,\mathrm{\mu m}$ to $y=44.87\,\mathrm{\mu m}$ in the $y$-direction. The leftmost $x$-coordinate of the experimental data is $x=0.65\,\mathrm{\mu m}$. The microwave amplitude at the position of the NV centers results from the interference between the microwave field generated by the spin waves and that from the stripline, and can be expressed as [21, 22, 26]
$B_{\mathrm{mw}}(x)$ (2)
$=\sqrt{B^{2}_{\mathrm{sw}}(x)+2B_{\mathrm{sw}}(x)B_{\mathrm{strip}}(x)\cos(k_{x}(x)x+\theta_{0})+B^{2}_{\mathrm{strip}}(x)},$
where $B_{\mathrm{strip}}(x)$ is the microwave amplitude distribution from the stripline, $B_{\mathrm{sw}}(x)$ is that from the spin waves, $k_{x}(x)$ is the spatial distribution of the $x$-component of the wavenumber and $\theta_{0}$ is the phase of the spin waves at the left edge of the field of view. When $B_{\mathrm{sw}}$ ($B_{\mathrm{strip}}$) is sufficiently larger than $B_{\mathrm{strip}}$ ($B_{\mathrm{sw}}$), $B_{\mathrm{sw}(\mathrm{strip})}\gg B_{\mathrm{strip}(\mathrm{sw})}$, the above equation can be approximated as
$B_{\mathrm{mw}}(x)\simeq B_{\mathrm{sw}(\mathrm{strip})}(x)+B_{\mathrm{strip}(\mathrm{sw})}(x)\cos(k_{x}(x)x+\theta_{0}).$ (3)
For YIG-A, under low input microwave power, the condition $B_{\mathrm{sw}}(x)>B_{\mathrm{mw}}(x)$ holds in the entire field of view (see Supplemental Materials for details [30]). Consequently, the mean line (the line passing through the center of the oscillations) of $B_{\mathrm{mw}}(x)$ shown in Fig. 1(d) corresponds to $B_{\mathrm{sw}}(x)$, while the oscillatory component corresponds to $B_{\mathrm{strip}}(x)$. From the decay of this mean line, we can estimate that the decay length of the spin waves is approximately $50\,\mathrm{\mu m}$.
III.2 From linear to nonlinear regime
We examine the case where the input power is varied in 10 steps from $5\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$. Figure 2(a) displays $B_{\mathrm{mw}}(x)$ at five selected power values (see Supplemental Materials for the full dataset [30]). The triangular markers with error bars represent the experimental data, while the solid black lines indicate the fitting results discussed later. The integrated region along the $y$-direction is the same as in Fig. 1(d). From $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to about $P_{\mathrm{mw}}=17\,\mathrm{dBm}$, the shape of the amplitude distribution remains unchanged; however, for input power above this range, significant modulation of the distribution is observed.
Figure 2: (a) One-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained at five different input microwave powers from $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ to $25\,\mathrm{dBm}$ for YIG-A. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data between large input microwave power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$) and small input microwave power ($P_{\mathrm{mw}}=11\,\mathrm{dBm}$) for YIG-A. (c) Microwave amplitude and (d) Wavenumber distribution from the spin waves for each input microwave power calculated from the fitting results for YIG-A.
Figure 2(b) compares $B_{\mathrm{mw}}(x)$ measured at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=25\,\mathrm{dBm}$. The first notable feature is that the shapes of the mean lines of $B_{\mathrm{mw}}(x)$ differ between the two power levels. This variation arises because $B_{\mathrm{strip}}(x)$ increases linearly with the input microwave power, whereas $B_{\mathrm{sw}}(x)$ exhibits a nonlinear response. According to Eq. (3), at low input powers, the mean line primarily reflects $B_{\mathrm{sw}}(x)$. In contrast, at high input powers, the mean line is dominated by the contribution from the stripline (see Supplemental Materials for details [30]). Furthermore, in the region of smaller $x$-coordinate [see the inset of Fig. 2(b)], the spacing between the peaks of the microwave amplitude oscillations is noticeably shorter at $P_{\mathrm{mw}}=25\,\mathrm{dBm}$ compared to $11\,\mathrm{dBm}$. This increase in the wavenumber is likely due to a modulation of the dispersion relation, caused by a reduction in the static magnetization along the $y$-direction resulting from the increase in the spin-wave amplitude [13].
Next, we quantitatively analyze the observed nonlinear dynamics by fitting the measured microwave amplitude distribution. Based on Eq. (2), we extract the microwave amplitude distribution from the spin waves $B_{\mathrm{sw}}(x)$, the wavenumber distribution of the spin waves $k_{x}(x)$, and the microwave amplitude distribution from the stripline $B_{\mathrm{strip}}(x)$. For the fitting, we assume the following expressions for $B_{\mathrm{sw}}(x)$, $k_{x}(x)$, and $B_{\mathrm{strip}}(x)$:
$\begin{split}B_{\mathrm{sw}}(x)&=B^{0}_{\mathrm{sw}}e^{-\quantity(\frac{x-x_{0}}{l_{\mathrm{sw}}})^{p_{\mathrm{sw}}}},\\ k_{x}(x)&=k_{0}+k_{1}e^{-\quantity(\frac{x-x_{0}}{l_{\mathrm{k}}})^{p_{\mathrm{k}}}},\\ B_{\mathrm{strip}}(x)&=B^{0}_{\mathrm{strip}}e^{-\quantity(\frac{x-x_{0}}{l_{\mathrm{strip}}})^{p_{\mathrm{strip}}}},\end{split}$ (4)
where $x_{0}$ denotes the coordinate of the left edge of the field of view ($x_{0}=0.65\,\mathrm{\mu m}$ in this measurement). For $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$, we adopt stretched exponential functions. The function for the spin waves (stripline) is characterized by the amplitude at $x=x_{0}$, $B^{0}_{\mathrm{sw}}=B_{\mathrm{sw}}(x_{0})$ ($B^{0}_{\mathrm{strip}}=B_{\mathrm{strip}}(x_{0})$), the decay length $l_{\mathrm{sw}}$ ($l_{\mathrm{strip}}$), and the stretch factor $p_{\mathrm{sw}}$ ($p_{\mathrm{strip}}$). In the linear regime, $B_{\mathrm{sw}}(x)$ generally decays exponentially; however, in the nonlinear regime, rapid decay of the amplitude near the stripline has been reported [13]. Therefore, we introduce a stretch factor into the exponential function to capture this behavior (see Supplemental Materials for details [30]). For the wavenumber distribution, we adopt a function consisting of a stretched exponential function corresponding to the spatial modulation of the wavenumber and a constant term $k_{0}$. The stretched exponential function is characterized by the wavenumber modulation at $x=x_{0}$, $k_{1}$, the decay length $l_{k}$, and the stretch factor $p_{\mathrm{k}}$.
When performing the fitting, we note that Eq. (2) is symmetric with respect to $B_{\mathrm{sw}}(x)$ and $B_{\mathrm{strip}}(x)$. Starting from random initial values makes it impossible to distinguish between these two contributions. To resolve this ambiguity, we use the microwave amplitude distribution obtained under the same input microwave power but with the direction of the external magnetic field reversed as a reference. When the magnetic field is reversed, the amplitude of the spin waves propagating in the +x+x-direction decreases due to a reduction of the spin-wave excitation efficiency of the stripline [21, 37]. Moreover, the microwave field from the spin waves propagating in the +x+x-direction in this configuration does not induce transitions between the NV spins’ $m_{s}=0$ and $m_{s}=-1$ states because of its polarization [21, 38]. Thus, we assume that the reference microwave amplitude distribution arises primarily from the stripline. Accordingly, we first fit the reference distribution and then use the obtained fitting parameters as the initial values for $B_{\mathrm{strip}}(x)$ in Eq. (4) (see Supplemental Materials for details [30]).
The fitted curves are shown in Fig. 2(a) as solid black lines. For each input microwave power, the fitted curve reproduces the experimental data well. In Fig. 2(c), we present the estimated microwave amplitude distribution from the spin waves $B_{\mathrm{sw}}(x)$ obtained from these fitting results. Here, we define the left end of the plot at $x=0.91\,\mathrm{\mu m}$, corresponding to the region where the amplitude variation from the stripline becomes relatively small. For input powers ranging from $P_{\mathrm{mw}}=5\,\mathrm{dBm}$ to approximately $19\,\mathrm{dBm}$, $B_{\mathrm{sw}}(x)$ increases with increasing input power and decays gradually over a similar decay length within the field of view. However, at microwave powers above this range, it no longer increases and instead decays rapidly near the stripline. In Fig. 2(d), we show the wavenumber distribution $k_{x}(x)$ estimated by fitting for each input microwave power. As the input power increases, the wavenumber also increases. While no significant spatial variation is observed within the field of view at lower input powers, noticeable spatial variations emerge from around $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. At the maximum power of $25\,\mathrm{dBm}$, the wavenumber varies by more than 20% across the field of view.
III.3 Nonlinear behavior of spin-wave amplitude
Now, we conduct a quantitative analysis of the spin-wave amplitude in the nonlinear regime. We convert the obtained microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$ into the corresponding spin-wave amplitude distribution. $B_{\mathrm{sw}}(x)$ is related to the spin-wave amplitude distribution in the $x$-direction $m_{x}(x)$ by the following equation [21, 22, 26]:
$B_{\mathrm{sw}}(x)=\frac{\mu_{0}}{2}e^{-k_{x}(x)z_{d}}(1-e^{-k_{x}(x)d})(1+\eta_{k_{x}(x)})m_{x}(x),\\$ (5)
where $\eta_{k_{x}(x)}$ is the ellipticity of the spin-wave precession, which can be obtained analytically from the experimental conditions (See Appendix B of Ref. [26] for details). Figure 3(a) shows the results of converting $B_{\mathrm{sw}}(x)$, shown in Fig. 2(a), into $m_{x}(x)$ using Eq. (5). As with the microwave amplitude distribution, the decay of the spin-wave amplitude near the stripline becomes more pronounced above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. The nonlinearity of the spin-wave amplitude is also reproduced by numerical simulations (see Supplemental Materials for details [30]), which validates our results.
Figure 3(b) displays a heatmap of $m_{x}(x)$ from Fig. 3(a) as a function of the input microwave amplitude $H_{\mathrm{mw}}\,(\coloneqq\sqrt{P_{\mathrm{mw}}})$. Up to approximately $H_{\mathrm{mw}}\sim 10\,\sqrt{\mathrm{mW}}$, the spin-wave amplitude around the left edge of the field of view (small $x$ region) increases with the increase of $H_{\mathrm{mw}}$, and gradually decays along the propagation direction. Beyond this range, however, the amplitude no longer increases and instead decays rapidly from the left edge, signaling peculiar nonlinearity.
Figure 3: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each input microwave power for YIG-A. (b) Heatmap of the spin-wave amplitude distribution as a function of the input microwave amplitude. (c)(d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=13\,\mathrm{\mu m}$.
Figure 3(c) shows the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ as a function of the input microwave amplitude. The red dashed line represents a linear fit through the origin based on data from the two smallest input powers ($P_{\mathrm{mw}}=5,11\,\mathrm{dBm}$). The spin-wave amplitude begins to deviate from linear behavior around $\mu_{0}m_{x}=20\,\mathrm{mT}$, increases up to approximately $30\,\mathrm{mT}$, and then begins to decay. This deviation from linearity at a certain microwave power is consistent with previous studies using VNA and BLS [12, 8, 9, 13]. Figure 3(d) shows the dependence of the spin-wave amplitude at $x=13\,\mathrm{\mu m}$. The amplitude deviates from linear behavior at the same $H_{\mathrm{mw}}$ as $x=4\,\mathrm{\mu m}$, saturates around $20\,\mathrm{mT}$, and then begins to decay.
III.4 Theoretical interpretation: four-magnon scattering
The observed nonlinear decay of the spin-wave amplitude can be attributed to multi-magnon scattering. In this process, excited magnons characterized by frequency $f_{\mathrm{sw}}$ and wavevector $\bm{k}$ are scattered into magnons with different frequencies and wavevectors while conserving the total energy and wavevector, and this scattering process becomes pronounced once the spin-wave amplitude exceeds a certain threshold. In this subsection, following Refs. [39, 40, 41, 42, 11, 2, 15], we first formulate the time evolution of magnon operators under the Hamiltonian that includes nonlinear terms and calculate the threshold magnon amplitude at which nonlinearity becomes significant. We then perform a quantitative comparison with the experimental results.
The magnon Hamiltonian $\hat{H}$, including nonlinear terms, can be expressed as [15, 2]
$\hat{H}=\hbar\sum_{\bm{k}}$ $\omega_{\bm{k}}\hat{c}_{\bm{k}}^{{\dagger}}\hat{c}_{\bm{k}}$ (6)
$+\sum_{\bm{k_{1}}+\bm{k_{2}}=\bm{k_{3}}+\bm{k_{4}}}H_{\bm{k_{1}},\bm{k_{2}},\bm{k_{3}},\bm{k_{4}}}\hat{c}_{\bm{k_{1}}}^{{\dagger}}\hat{c}_{\bm{k_{2}}}^{{\dagger}}\hat{c}_{\bm{k_{3}}}\hat{c}_{\bm{k_{4}}},$
where the first term represents the linear term describing the dispersion relation of the spin waves, while the second term corresponds to the nonlinear four-magnon scattering process characterized by the coupling strength $H_{\bm{k_{1}},\bm{k_{2}},\bm{k_{3}},\bm{k_{4}}}$ (see Supplemental Materials for the explicit expressions for each term and their derivations [30]). Although nonlinear processes such as three-magnon scattering also appear among the nonlinear terms, they do not contribute due to conservation laws governing frequency and wavevector. Therefore, we consider only the four-magnon scattering term described above (see Supplemental Materials for details [30]).
In the Heisenberg picture, the time evolution of the magnon operator $\hat{c}_{\bm{k}}$ follows the Heisenberg equation:
$i\hbar\frac{\mathrm{d}}{\mathrm{d}t}\hat{c}_{\bm{k}}=\quantity[\hat{c}_{\bm{k}},\hat{H}$]. (7)
Here, we consider the situation where only magnons with wavevector $\bm{k}$ are excited by a microwave field. These excited magnons subsequently undergo four-magnon scattering such that $\bm{k_{1}}=\bm{k}+\bm{q}$, $\bm{k_{2}}=\bm{k}-\bm{q}$, and $\bm{k_{3}}=\bm{k_{4}}=\bm{k}$ in Eq. (6) [42, 11]. In the rotating frame defined as $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\,(=\hat{c}_{\bm{k}-\bm{q}}e^{i\omega_{\bm{k}-\bm{q}}t})$, the time evolution of the magnon annihilation operator obeys the following differential equation:
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}=-\Gamma_{\bm{k}-\bm{q}}\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}-iH_{\bm{k}-\bm{q},\bm{k}+\bm{q},\bm{k},\bm{k}}\hat{c}_{\bm{k}+\bm{q}}^{{}^{\prime}{\dagger}}\hat{c}^{{}^{\prime}}_{\bm{k}}\hat{c}^{{}^{\prime}}_{\bm{k}},$ (8)
where we have introduced the decay term $\Gamma_{\bm{k}-\bm{q}}$. Following the same discussion, the time evolution of $\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}+\bm{q}}$ obeys the following differential equation:
$\frac{\mathrm{d}}{\mathrm{d}t}\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}+\bm{q}}=-\Gamma_{\bm{k}+\bm{q}}\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}+\bm{q}}+iH^{*}_{\bm{k}+\bm{q},\bm{k}-\bm{q},\bm{k},\bm{k}}\hat{c}_{\bm{k}-\bm{q}}^{{}^{\prime}}\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}}\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}}.$ (9)
For the nonlinear term in Eqs. (8) and (9), we assume that $\hat{c}^{{}^{\prime}}_{\bm{k}}$ attains the steady-state value $\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}$ under resonant microwave excitation [43]. By combining Eqs. (8) and (9), we derive the following differential equation for $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}$:
$\quantity(\frac{\mathrm{d}}{\mathrm{d}t}+\Gamma_{\bm{k}+\bm{q}})$ $\quantity(\frac{\mathrm{d}}{\mathrm{d}t}+\Gamma_{\bm{k}-\bm{q}})\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}$ (10)
$=H^{*}_{\bm{k}+\bm{q},\bm{k}-\bm{q},\bm{k},\bm{k}}H_{\bm{k}-\bm{q},\bm{k}+\bm{q},\bm{k},\bm{k}}\quantity|\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}|^{4}\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}.$
Now, we define the magnon generation rate $\gamma_{\bm{k}-\bm{q}}$ by assuming that $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\propto e^{\gamma_{\bm{k}-\bm{q}}t}$. Taking the limit $\bm{q}\rightarrow 0$ in Eq. (10), we obtain
$\gamma_{\bm{k}}=-\Gamma_{\bm{k}}+\quantity|H_{\bm{k},\bm{k},\bm{k},\bm{k}}|\quantity|\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}|^{2}.$ (11)
This equation indicates that the generation rate of magnons produced by scattering changes from negative to positive as the magnon amplitude increases from zero. The resulting sign change corresponds to a threshold at which the excited coherent spin waves begin to decay via four-magnon scattering. At this threshold, the magnon amplitude is given by
$\quantity|\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}|=\sqrt{\frac{\Gamma_{\bm{k}}}{\quantity|H_{\bm{k},\bm{k},\bm{k},\bm{k}}|}}.$ (12)
By introducing a dimensionless parameter $\lambda_{\bm{k}}$ to characterize the four-magnon scattering, $H_{\bm{k},\bm{k},\bm{k},\bm{k}}$ is expressed as
$H_{\bm{k},\bm{k},\bm{k},\bm{k}}=\lambda_{\bm{k}}\frac{\hbar\gamma^{\prime 2}_{e}}{V},$ (13)
where $V$ is the volume of the magnetic film and $\gamma^{\prime}_{e}=2\pi\gamma_{e}$. By utilizing the expression above, the spin-wave amplitude in the $x$-direction at the threshold $m^{\mathrm{th}}_{x,\bm{k}}$ can be calculated as
$m^{\mathrm{th}}_{x,\bm{k}}=(u_{\bm{k}}+v_{\bm{k}})\sqrt{\frac{2\Gamma_{\bm{k}}}{\quantity|\lambda_{\bm{k}}|\gamma^{\prime}_{e}M_{\mathrm{s}}}}M_{\mathrm{s}},$ (14)
where $u_{\bm{k}}$ and $v_{\bm{k}}$ are dimensionless parameters of the Bogoliubov transformation [41] (see Supplemental Materials for details [30]), and $M_{\mathrm{s}}$ is the saturation magnetization of the magnetic film.
We calculate the spin-wave amplitude threshold corresponding to the experimental results using this expression. Table 1 summarizes the relevant experimental parameters. For YIG-A, we use the parameters listed in the left column of Table 1. The saturation magnetization is taken as $M_{\mathrm{s}}=141\,\mathrm{mT}$, as measured by SQUID (superconducting quantum interference device) magnetometry at 300 K. The anisotropy field $H_{\mathrm{ani}}$ is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=169.6\,\mathrm{mT}$ and the saturation magnetization as $H_{\mathrm{ani}}=M_{s}-M_{\mathrm{eff}}$. The decay rate $\Gamma_{\bm{k}}$ is obtained from the spin-wave lifetime, which is estimated by dividing the decay length $l_{d}=54.21\,\mathrm{\mu m}$ (determined from fitting the data at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$) by the group velocity of the spin waves. The four-magnon scattering coefficient $\lambda_{\bm{k}}$ and the Bogoliubov transformation parameters $u_{\bm{k}}$ and $v_{\bm{k}}$ are calculated from analytical expressions using the experimental parameters (see Supplemental Materials for the explicit expressions [30]). Based on these parameters, the spin-wave amplitude threshold for the instability due to four-magnon scattering is calculated to be $\mu_{0}m^{\mathrm{th}}_{x,\bm{k}}=17.8\,\mathrm{mT}$.
It is important to examine how the above discussion applies to our experiment. The black dotted lines in Figs. 3(c) and (d) indicate the threshold amplitude obtained in this way. In particular, regarding Fig. 3(c), which displays the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ near the stripline, the threshold for the emergence of nonlinearity and the experimental results agree satisfactorily. This result demonstrates that the nonlinear dynamics can be quantitatively visualized using NV centers. The ability to experimentally detect and quantitatively characterize the transition from the linear to the nonlinear regime, consistent with theoretical predictions, constitutes the central achievement of the present work.
YIG-A YIG-B
Film thickness $d$ $54\,\mathrm{nm}$ $109\,\mathrm{nm}$
External magnetic field $B_{0}$ $5.17\,\mathrm{mT} 22$$2.45\,\mathrm{mT}$
Saturation magnetization $M_{\mathrm{s}}$ $141\,\mathrm{mT}$ $162\,\mathrm{mT}$
Anisotropy field $H_{\mathrm{ani}}$ $28.6\,\mathrm{mT}$ $15.6\,\mathrm{mT}$
Spin-wave frequency $f_{\mathrm{sw}}$ $166\,\mathrm{MHz} 22$$243.1\,\mathrm{MHz}$
Spin-wave wavenumber $k_{x}$ $1.45\,\mathrm{rad/\mu m}$ $1.27\,\mathrm{rad/\mu m}$
Decay rate $\Gamma_{\bm{k}}$ $12.6\,\mathrm{MHz}$ $4.78\,\mathrm{MHz}$
Four-magnon scattering coefficient $\lambda_{\bm{k}}$ -0.155 -0.173
Bogoliucdbov transformation parameter $u_{\bm{k}}$ 1.11 1.10
Bogoliubov transformation parameter $v_{\bm{k}}$ 0.472 0.449
Table 1: Experimental parameters of YIG-A and YIG-B.
III.5 Magnetization reduction in the nonlinear regime
Having examined the spin-wave amplitude, we next analyze the spatial distribution of the wavenumber. The observed increase in the wavenumber reflects a decrease in the static magnetization along the $y$-direction (i.e., the direction of the saturation magnetization) $M_{y,s}$ and the corresponding effective magnetization $M_{\mathrm{eff}}$ due to an increase in the spin-wave amplitude [Fig. 4(a)] [13, 14]. To calculate the effective magnetization, we utilize the dispersion relation of surface spin waves [Eq. (1)]. At each position, the effective magnetization $M_{\mathrm{eff}}(x)$ corresponding to the spin-wave frequency $f_{\mathrm{sw}}=2165.8\,\mathrm{MHz}$ is calculated from the wavenumber $k_{x}(x)$. As the dispersion relation holds for a small spin-wave amplitude, there may be deviations from Eq. (1) for large amplitudes, particularly in nonlinear regions. However, the coherent spin-wave amplitude with the excitation microwave frequency decays rapidly near the stripline in the nonlinear region, and the amplitude is small compared to the saturation magnetization across most of the field of view. Therefore, we utilize the dispersion relation of the surface spin waves in Eq. (1) to derive the static magnetization distribution approximately.
Figure 4(c) plots the resulting spatial distribution of the static magnetization $M_{y,\mathrm{s}}(x)$. We convert the effective magnetization distribution $M_{\mathrm{eff}}(x)$ into the static magnetization distribution $M_{\mathrm{y,s}}(x)$ by removing the contribution from the surface anisotropy $\quantity(M_{\mathrm{y,s}}(x)=M_{\mathrm{eff}}(x)+H_{\mathrm{ani}})$. Figure 4(c) shows that, in regions with small input microwave power, though there is a uniform reduction as the power increases, the static magnetization remains almost constant within the field of view. However, in addition to the uniform reduction, significant spatial variation within the field of view appears above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$, and at the maximum input power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$), a variation of about $6\,\mathrm{mT}$ within the field of view is observed.
Figure 4: (a)(b) Schematic of the reduction of the static magnetization along the $y$-axis, (a) considering contributions from all spin waves. (b) considering only contributions from excited coherent spin waves. (c) Static magnetization distribution obtained by converting the wavenumber distribution for YIG-A. (d) Estimation of the static magnetization distribution considering only contributions from the excited coherent spin waves for YIG-A.
It is instructive to examine how the surface spin waves contribute to the reduction in static magnetization [Fig. 4(b)]. When considering only the contribution from the coherent surface spin waves with the excitation frequency, the magnetization in the $y$-direction at a specific position and time is expressed as
[…]
where $m_{x}(t)$ and $m_{z}(t)$ are the precession of the magnetization components along the $x$- and $z$-directions of the surface spin waves, respectively:
[…]
Consequently, the averaged static magnetization $M^{{}^{\prime}}_{y,\mathrm{s}}$, which accounts only for the contribution from the coherent surface spin waves, can be obtained by averaging the phase of the spin waves $\theta\,(=2\pi f_{\mathrm{sw}}t)$ as
$M^{{}^{\prime}}_{y,\mathrm{s}}=\frac{1}{2\pi}\int_{0}^{2\pi}\sqrt{M^{2}_{s}-\quantity(m_{x}\cos\theta)^{2}-\quantity(\eta_{k}m_{x}\sin\theta)^{2}}\,d\theta.$ (17)
Figure 4(d) shows the effective static magnetization distribution $M^{{}^{\prime}}_{y,\mathrm{s}}(x)$, which is calculated from the spin-wave amplitude distribution in Fig. 3(a) using Eq. (17). The left edge of the plot is defined at $x=0.91\,\mathrm{\mu m}$. At small input microwave powers in the linear regime, $M^{{}^{\prime}}_{y,\mathrm{s}}$ exhibits a slight decrease only near the left edge of the field of view ($x\lesssim 10$ $\mathrm{\mu m}$), which is in clear contrast to the observation that the static magnetization derived from the wavenumber $M_{\mathrm{y,s}}(x)$ shown in Fig. 4(c)] shows a prominent decrease across the entire field of view. These results imply that the decay of the surface spin waves is not solely due to relaxation into the lattice but also involves scattering by other magnons, such as thermal magnons. In other words, the excited surface spin waves are transformed into other modes, leading to a persistent reduction in the static magnetization across the field of view [41].
Around maximum input microwave powers in the nonlinear regime, the reduction in static magnetization attributable to surface spin waves becomes almost negligible at a distance of about $5\,\mathrm{\mu m}$ from the stripline. In contrast, the static magnetization estimated from the wavenumber [Fig. 4(c)] continues to change gradually across the field of view. This observation suggests that most of the modulation in magnetization is caused by additional magnons generated through four-magnon scattering from the excited surface spin waves.
As a note, we investigate the possible role of Joule heating resulting from the application of a large-amplitude microwave, which might cause the sample to heat up. Such heating could potentially create a temperature gradient along the $x$-direction, thereby causing a spatial variation in the saturation magnetization and wavenumber. However, we can safely exclude this possibility. We perform similar measurements at a different magnetic field, where the spin-wave excitation is more inefficient and the spin waves are less likely to reach the nonlinear regime, and confirm that the wavenumber along the $x$-direction remains almost constant even at high microwave power (see Supplemental Materials for details [30]).
III.6 Comparison with a different YIG film
So far, we have quantitatively imaged the amplitude and wavenumber distributions of spin waves in YIG-A over a wide range of input microwave powers. In particular, we have confirmed that the threshold of nonlinearity coincides with the threshold determined by four-magnon scattering. As shown in Eq. (14), this threshold depends on the relaxation rate of the spin waves. It is therefore important to investigate how our interpretation can be generalized. One possible approach to demonstrate the validity of our results is to check whether the amplitude indirectly inferred in Ref. [13] is consistent with theory. However, a straightforward and quantitative comparison is difficult for two reasons. First, the present work investigates a two-dimensional extended YIG thin film, whereas Ref. [13] studies a waveguide. As a result, key parameters of the magnon Hamiltonian, including dipolar interactions, differ substantially, making it difficult to apply the theory used in our paper as is. Second, in Ref. [13], the spin-wave precession angle is inferred from a numerically computed dispersion by imposing a uniform tilt $\theta$ and then allowing the spins to precess. In this procedure, however, the ellipticity of spin-wave precession is not taken into account, and it is unclear how faithfully the imposed $\theta$ maps onto the actual dynamic magnetization amplitude in the experimental system. For these reasons, the indirectly determined threshold reported in Ref. [13] cannot be assessed in a rigorous manner. Instead, to substantiate the validity of both the theory and the measurements, we present results of YIG-B, which has a different relaxation rate.
The external magnetic field $B_{0}$ is set to $22.45\,\mathrm{mT}$ when measuring YIG-B. The resonance frequency of the NV spins $f_{\mathrm{NV}}$ is $2243.1\,\mathrm{MHz}$, and the wavenumber of the surface spin waves at this frequency is around $1.3\,\mathrm{rad/\mu m}$ (see Supplemental Materials for details [30]).
We sweep the input microwave power $P_{\mathrm{mw}}$ in 23 steps from $0\,\mathrm{dBm}$ to $22\,\mathrm{dBm}$. Figure 5(a) displays the microwave amplitude distribution $B_{\mathrm{mw}}(x)$ for five representative input powers (see Supplemental Materials for additional data [30]). The triangular markers with error bars indicate the experimental data. The integration region along the $y$-direction is the same as that used for YIG-A [see Figs. 1(d) and 2(a)]. Focusing on the data at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$, we observe that, compared with YIG-A, the decay of $B_{\mathrm{mw}}(x)$ from the left to the right end is small, reflecting a lower relaxation rate of the spin waves. Overall, the mean line of $B_{\mathrm{mw}}(x)$ attenuates by approximately 20% across the entire field of view, which corresponds to a decay length of about $300\,\mathrm{\mu m}$.
Figure 5: (a) Input microwave power dependence of the one-dimensional microwave amplitude distribution $B_{\mathrm{mw}}(x)$ obtained for YIG-B. The triangular markers with error bars represent experimental data, and the solid black lines correspond to the fitting results. (b) Comparison of $B_{\mathrm{mw}}(x)$ for data with large input microwave power ($P_{\mathrm{mw}}=22\,\mathrm{dBm}$) and data with small input microwave power ($P_{\mathrm{mw}}=6\,\mathrm{dBm}$) for YIG-B. (c) Microwave amplitude and (d) Wavenumber distribution from the spin waves for each input microwave power calculated from the fitting results for YIG-B.
In Fig. 5(b), $B_{\mathrm{mw}}(x)$ at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$ and $P_{\mathrm{mw}}=22\,\mathrm{dBm}$ are overlaid for direct comparison [see Fig. 2(b) for YIG-A]. Although there is no pronounced decay at $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the amplitude of the mean line is reduced by approximately 50% from the left to the right end of the field of view, indicating the emergence of nonlinear decay at this input power. Furthermore, when focusing on the smaller $x$-coordinate shown in the inset, a slight difference in wavenumber between the two input microwave powers is observed. However, the modulation is less significant than that seen in YIG-A.
We follow the same procedure as YIG-A to fit $B_{\mathrm{mw}}(x)$, with two differences. First, for YIG-B, across all regions and at every input microwave power, the reference microwave distribution is consistently smaller than the measured microwave amplitude distribution (see Supplemental Materials for details [30]). Consequently, we employ Eq. (3) for the fitting. Second, because the observed wavenumber modulation is small, we adopt the following linear approximation for the wavenumber distribution to describe the minor spatial modulation:
$k_{x}(x)=k_{0}-k_{1}\frac{(x-x_{0})}{l_{\mathrm{k}}}.$ (18)
For each input microwave power, the fitted curve reproduces the experimental data precisely, as shown by the solid black lines in Fig. 5(a).
In Fig. 5(c), we plot the microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$, estimated from the fitting results at eleven different input microwave powers. The left end of the plot is set at $x=0.91\,\mathrm{\mu m}$. Notably, the amplitude at the left end continues to increase up to the maximum input power of $P_{\mathrm{mw}}=22\,\mathrm{dBm}$. In contrast, the amplitude at the right end saturates at around $B_{\mathrm{sw}}=0.5\,\mathrm{mT}$, indicating the onset of nonlinearity. Figure 5(d) shows the estimated wavenumber distribution $k_{x}(x)$ for each input microwave power. The modulation of $k_{x}(x)$ across the field of view is minimal; even at the maximum input microwave power of $P_{\mathrm{mw}}=22\,\mathrm{dBm}$, the variation from the left to the right end is less than $0.1\,\mathrm{rad/\mu m}$. These observations are markedly different from those for YIG-A, as shown in Figs. 2(c) and (d).
Figure 6: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each input microwave power for YIG-B. (b) The heatmap of the spin-wave amplitude distribution for each input microwave amplitude for YIG-B. (c) (d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=78\,\mathrm{\mu m}$ for YIG-B.
As we did for YIG-A, we use Eq. (5) to convert $B_{\mathrm{sw}}(x)$, shown in Fig. 5(c), into the corresponding spin-wave amplitude distributions $m_{x}(x)$. Figure 6(a) shows the converted spin-wave amplitude, and Fig. 6(b) displays a heatmap of $m_{x}(x)$ for all data, together with the individual profiles from Fig. 6(a). To investigate the nonlinearity of the spin waves, we examine the dependence of the spin-wave amplitude on the input microwave amplitude at two positions: $x=4\,\mathrm{\mu m}$ on the left side of the field of view and $x=78\,\mathrm{\mu m}$ on the right side. Figure 6(c) shows the dependence of the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ on the input microwave amplitude. The red dashed line represents a linear fit through the origin, obtained using data from the five smallest input microwave amplitudes (up to $P_{\mathrm{mw}}=10\,\mathrm{dBm}$). The data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$. Although the amplitude does not immediately saturate, it deviates from the linear trend and the gradient continues to decrease. Similarly, Fig. 6(d) shows the dependence of the spin-wave amplitude at $x=78\,\mathrm{\mu m}$. As for the data at $x=4\,\mathrm{\mu m}$, nonlinearity appears when the amplitude reaches around $10\,\mathrm{mT}$. However, in contrast to the left side, the amplitude at $x=78\,\mathrm{\mu m}$ does not increase significantly in the nonlinear region and appears to saturate around $13\,\mathrm{mT}$. These behaviors are qualitatively similar to those observed for YIG-A, but the threshold at which nonlinearity emerges is smaller than in YIG-A.
We compare the obtained results with the theoretical threshold given by Eq. (14). The right column of Table 1 lists the parameters used for the calculation. The saturation magnetization is taken as $M_{\mathrm{s}}=162\,\mathrm{mT}$, as measured by SQUID magnetometry. The anisotropy field is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=177.6\,\mathrm{mT}$ and the saturation magnetization. The decay rate is calculated from the decay length $l_{d}=318\,\mathrm{\mu m}$, obtained from the fitting result at $P_{\mathrm{mw}}=6\,\mathrm{dBm}$. With these parameters, the threshold spin-wave amplitude for the onset of four-magnon decay is estimated to be $\mu_{0}m^{\mathrm{th}}_{x}=10.5\,\mathrm{mT}$. As expected, this threshold is lower than that of YIG-A, reflecting the smaller decay rate in the present case. As shown in Figs. 6(c) and (d), the experimental data indicate that nonlinearity begins to occur when the spin-wave amplitude exceeds approximately $10\,\mathrm{mT}$ (marked by the horizontal lines in the graphs), again in good agreement with this theoretical prediction.
Unlike in YIG-A, the absence of significant wavenumber modulation in YIG-B can be attributed to the smaller spin-wave amplitude, even at maximum power. As shown in Fig. 3(c), at $x=4\,\mathrm{\mu m}$, the extrapolated value from the linear fit of the spin-wave amplitude is approximately $70\,\mathrm{mT}$ at the maximum excitation microwave amplitude, whereas it is around $40\,\mathrm{mT}$ for YIG-B. Consequently, even at maximum power, the lower amplitude of the excited spin waves in YIG-B results in only minor variations in the static magnetization across the field of view.
IV Conclusion and Perspectives
To conclude, we have employed QDM to investigate the nonlinear dynamics of surface spin waves quantitatively. For YIG-A, we swept the input microwave power to map the microwave amplitude distribution, revealing both nonlinear behavior in the amplitude profile and shifts in the wavenumber resulting from a reduction in the saturation magnetization. Quantitative analysis of these data showed that nonlinearity in the amplitude emerges once the spin-wave amplitude exceeds a specific threshold, consistent with theoretical predictions based on four-magnon scattering. Furthermore, as the spin-wave amplitude increases, we observed modulation of the wavenumber within the field of view, and our quantitative analysis revealed that this modulation largely originates from the spin waves generated by multi-magnon scattering. Similar measurements conducted on YIG-B, which exhibits a different relaxation rate, confirmed that the threshold for the onset of nonlinearity also agrees with the theoretical predictions.
These results deepen our understanding of nonlinear spin-wave dynamics and underscore the utility of NV centers as sensors for the quantitative investigation of nonlinear spin-wave phenomena, which have previously been accessible only at a qualitative level.
As a future extension of this work, one could consider the quantitative visualization of the incoherent magnons generated by four-magnon scattering. NV centers can quantitatively measure the magnetic noise at their resonance frequency via longitudinal relaxation measurements. Therefore, by exciting surface spin waves with frequencies detuned from the NV spin resonance and monitoring the relaxation rate of the NV spins while varying the input microwave power, it becomes possible to extract quantitative information about the incoherent magnons produced by nonlinear scattering [37]. Furthermore, under the same conditions, quantitative detection of the coherent spin waves detuned from the NV resonance frequency is possible by employing sensing based on the AC Zeeman effect [26].
In addition, nonlinear phenomena regarding spin waves are not limited to surface spin waves. For example, parametric excitation, where spin waves are generated by microwave fields at twice the frequency of the spin waves oriented along the saturation magnetization direction [3, 4, 5], is an important example. Furthermore, in the case of large-amplitude spin waves, the propagation of spin-wave bullets, which behave as solitons [44], represents another notable nonlinear effect. NV centers are expected to serve as powerful probes for investigating these phenomena.
The nonlinear dynamics of spin waves are also anticipated to have applications in spintronics devices [45, 46]. Recent developments of magnonics has been especially active in spin waves in the shorter-wavelength, higher-frequency regime [47], within which nonlinearity of the spin waves in particular holds substantial potential for efficient spin-wave excitation [48] and applications for logic devices [49] and neuromorphic computing [50]. Methods capable of quantitatively evaluating these operations can provide valuable guidelines for device design.
Acknowledgements
This work was partially supported by JST, CREST Grant No. JPMJCR23I2, Japan; Grants-in-Aid for Scientific Research (Nos. JP22K03524, JP22KJ1058, JP22KJ1059, JP23K25800, JP23K21077, JP24K21194, JP24H01666 and JP25H01248); the Mitsubishi Foundation (Grant No. 202310021); Seiko Instruments; Advanced Technology Foundation Research Grants; the Cooperative Research Project of RIEC, Tohoku University; “Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” (No. JPMXP1222UT1131) of the Ministry of Education, Culture, Sports, Science and Technology of Japan (MEXT). K.O. and M.T. acknowledge supports from FoPM, WINGS Program, The University of Tokyo.